OpenStudy (anonymous):

Examine the formulas below for a $5000 investment. Determine the number of interest periods, the nominal rate and the effective rate. a. A=5000(1+ (0.0345/4))^4t Nominal rate:_____% Effective rate:_____% b. A=(1+(0.052/12))^12t Nominal rate:______% Effective rate:______% c. A= (1+0.36)^t Nominal rate:_____% Effective rate:______% d. A=e^0.0641t Nominal rate:_____% effective rate:_______%

OpenStudy (goformit100):

d. A=e^0.0641t Nominal rate:_____% effective rate:_______%

OpenStudy (anonymous):

yes thats the last problem...how can we solve it tho?

OpenStudy (mertsj):

I don't know those terms. I'll have to look them up or maybe someone will come along who knows them.

OpenStudy (anonymous):

hmmm the nomical and effective rate you mean?

OpenStudy (mertsj):

Yes. The nominal rate is the rate without compounding.

OpenStudy (anonymous):

hmmm im still a little lost...

OpenStudy (mertsj):

So for the first one, I think that would be the 3.45%

OpenStudy (anonymous):

so i have... a.) N.R: 3.45% E.R: b.)N.R: 5.2 E.R: c.) N.R: E.R: d.) N.R: 6.41 E.R: im missing some

terenzreignz (terenzreignz):

P = principal, \[\huge P\left(1+\frac r {n}\right)^{nt}\]

OpenStudy (anonymous):

so...

terenzreignz (terenzreignz):

Well, nominal is r, that's certain...

jimthompson5910 (jim_thompson5910):

In the formula A = P(1+r/n)^(n*t) the nominal interest rate is r the effective interest rate (E) is found by this formula E = (1 + r/n)^(n*t) - 1

OpenStudy (mertsj):

Now the effective rate... I have found that the effective rate is: